# STEP: Simultaneous Tracking and Estimation of Pose for Animals and Humans

### v1

@@ Abstract
We introduce STEP, a novel framework utilizing Transformer-based discriminative model prediction for simultaneous tracking and estimation of pose across diverse animal species and humans. We are inspired by the fact that the human brain exploits spatiotemporal continuity and performs concurrent localization and pose estimation despite the specialization of brain areas for form and motion processing. Traditional discriminative models typically require predefined target states for determining model weights, a challenge we address through Gaussian Map Soft Prediction (GMSP) and Offset Map Regression Adapter (OMRA) Modules. These modules remove the necessity of keypoint target states as input, streamlining the process. Our method starts with a known target state initialized through a pre-trained detector or manual initialization in the initial frame of a given video sequence. It then seamlessly tracks the target and estimates keypoints of anatomical importance as output for subsequent frames. Unlike prevalent top-down pose estimation methods, our approach doesn’t rely on per-frame target detections due to its tracking capability. This facilitates a significant advancement in inference efficiency and potential applications. We train and validate our approach on datasets encompassing diverse species. Our experiments demonstrate superior results compared to existing methods, opening doors to various applications, including but not limited to action recognition and behavioral analysis.
@@ 1 Introduction
  • We conduct a comprehensive study on the pose estimation task within a discriminative model prediction framework. We utilize a Transformer-based model prediction module while simultaneously tracking the object of interest— an unexplored area in previous research.
@@ 3.1 Background
Figure 3: For a video sequence, columns display target score maps $\hat{K}_{\text{gm}}$ when initiated with bounding boxes, as depicted in (a) and (b). In (c), our STEP framework runs concurrently for each bounding box to track and estimate the pose of respective targets. Notably, observe the robustness of STEP in estimating keypoints, particularly in scenarios involving occlusion. A failure case is showcased in the last frame of (c) for the red bounded target, stemming from ambiguous activations within the red target’s score map, as evident in the last frame of (a). Additional qualitative results across various species are provided in the Supplementary material.
@@ 4.2.1 Automatic Initialization
Our proposed approach requires initialization with a bounding box to establish information on the target of interest. We conduct an experiment to assess the effect of initialization by employing a pre-trained Detector [55] to identify humans in the validation split of the CrowdPose Dataset [24]. The detector provides class-specific bounding boxes, and since CrowdPose is a human dataset, we filter out the predicted bounding boxes specific to the person category. However, the detector can sometimes mistakenly identify non-person objects as human or detect instances whose corresponding keypoint annotations are unavailable in the ground truth, which can hinder metric evaluation. To address this, we filter the predicted bounding boxes by selecting those with the maximum Intersection over Union (IoU) with the annotated boxes in the dataset. We run STEP on validation splits with initializations from the detector and with annotation available in the CrowdPose dataset and observe the minimal impact on metrics, as evident from Table 2. Initializing STEP with a pre-trained detector [55] has minimal impact on performance while fully automating the process.
@@ 4.3 Qualitative Results
A supplementary video is available, demonstrating the results on video sequences of animals from [50], humans from [11], and mice and fishes from [21]. We present a frame from the sequences in the supplementary video in Figure 5.
@@ 4.4 Inference Speed Analysis
The video also contains results using the top-down approach with STEP*, utilizing per-frame bounding box information. Since the detections for each frame are known in advance, STEP* can manage newly entered objects, maintain tracking, and estimate their respective poses.

### v2

@@ Abstract
We introduce STEP, a novel framework utilizing Transformer-based discriminative model prediction for simultaneous tracking and estimation of pose across diverse animal species and humans. We are inspired by the fact that the human brain exploits spatiotemporal continuity and performs concurrent localization and pose estimation despite the specialization of brain areas for form and motion processing. Traditional discriminative models typically require predefined target states for determining model weights, a challenge we address through Gaussian Map Soft Prediction (GMSP) and Offset Map Regression Adapter (OMRA) Modules. These modules remove the necessity of keypoint target states as input, streamlining the process. Our method starts with a known target state in the initial frame of a given video sequence. It then seamlessly tracks the target and estimates keypoints of anatomical importance as output for subsequent frames. Unlike prevalent top-down pose estimation methods, our approach doesn’t rely on per-frame target detections due to its tracking capability. This facilitates a significant advancement in inference efficiency and potential applications. We train and validate our approach on datasets encompassing diverse species. Our experiments demonstrate superior results compared to existing methods, opening doors to various applications, including but not limited to action recognition and behavioral analysis.
@@ 1 Introduction
  • We conduct a comprehensive study on pose estimation task within a discriminative model prediction framework. We utilize a Transformer-based model prediction module while simultaneously tracking the object of interest— an unexplored area in previous research.
@@ 3.1 Background
Figure 3: For a video sequence, columns display target score maps $\hat{K}_{\text{gm}}$ when initiated with bounding boxes, as depicted in (a) and (b). In (c), our STEP framework runs concurrently for each bounding box to track and estimate the pose of respective targets. Notably, observe the robustness of STEP in estimating keypoints, particularly in scenarios involving occlusion. A failure case is showcased in the last frame of (c) for the red bounded target, stemming from ambiguous activations within the red target’s score map, as evident in the last frame of (a).
@@ 4.2.1 Automatic Initialization
Our proposed approach requires initialization with a bounding box to establish information on the target of interest. We conduct an experiment to assess the effect of initialization by employing a pre-trained Detector [56] to identify humans in the validation split of the CrowdPose Dataset [25]. The detector provides class-specific bounding boxes, and since CrowdPose is a human dataset, we filter out the predicted bounding boxes specific to the person category. However, the detector can sometimes mistakenly identify non-person objects as human or detect instances whose corresponding keypoint annotations are unavailable in the ground truth, which can hinder metric evaluation. To address this, we filter the predicted bounding boxes by selecting those with the maximum Intersection over Union (IoU) with the annotated boxes in the dataset. We run STEP on validation splits with initializations from the detector and with annotation available in the CrowdPose dataset and observe the minimal impact on metrics, as evident from Table 2. Initializing STEP with a pre-trained detector [56] has minimal impact on performance while fully automating the process. Furthermore, our work addresses the fundamental challenge of simultaneous tracking and pose estimation in wild settings. We highlight that in constrained environments, such as controlled laboratories or zoo enclosures, animal localization is relatively straightforward, as most or all elements in the scene remain static. Simple bottom-up approaches can efficiently handle pose estimation and localization tasks [17], which can be easily integrated into our method and can be utilized as an initialization step.
@@ 4.3 Qualitative Results
A supplementary video is available (Link) demonstrating the results on video sequences of animals from [51], humans from [11], and mice and fishes from [22]. We present a frame from the sequences in the supplementary video in Figure 5. The video also contains results using the top-down approach with STEP*, utilizing per-frame bounding box information. Since the detections for each frame are known in advance, STEP* can manage newly entered objects, maintain tracking, and estimate their respective poses.
@@ 5 Conclusion
Acknowledgments
We thank Prof. Kazunori Yamada from Osaka University and Hisami-san from the Awaji Island Monkey Center for granting us access to the center’s video feeds for our analysis. We also thank Dr. Mami Matsuda for facilitating communication with the Awaji Monkey Center. Additionally, we acknowledge the support of the Jibaben Patel Chair in AI for this work.
Data Availability
All experiments are performed on publically available datasets: APT36K [51] (https://github.com/pandorgan/APT-36K), APT10K [54] (https://github.com/AlexTheBad/AP-10K), CrowdPose [25] (https://github.com/jeffffffli/CrowdPose) and TriMouse, Marmoset and Fish from DeepLabCut benchmark dataset [22] https://benchmark.deeplabcut.org/datasets.html. For Figure 2 and Figure 8(a), we used the video stream available at (https://www.youtube.com/live/lsxYH2XQQCg). We have made the STEP project page publicly accessible at Link, which includes access to supplementary video (https://youtu.be/TxSk3OtJ6R8) and code (https://github.com/shash29-dev/STEP).